OpenStudy (anonymous):

A boat travels at a rate of 15km/hr in still water . It travels 60km upstream in the same time that it travels 90km downstream. What is the rate of the current ?

OpenStudy (mathstudent55):

Current speed = c Boat speed is still water = s Boat speed upstream = s - c Boat speed downstream = s + c

OpenStudy (mathstudent55):

speed = distance/time

OpenStudy (mathstudent55):

distance = speed * time

OpenStudy (mathstudent55):

You understand the table so far?